# Rethinking Softmax: Self-Attention with Polynomial Activations

### v1

@@ A.1 Theoretical analysis
$\mathop{\mathbb{E}}(c_{11}^{2})$ $\approx\mathop{\mathbb{E}}(\sum_{\{i_{1},\dots,i_{p}\}\in\{1,\dots,D\}}\prod_{k=1}^{p}a_{1,i_{k}}^{2}b_{i_{k},1}^{2})$ (A.30)
@@
$\mathop{\mathbb{E}}(\|\frac{\partial XQK^{T}X^{T}}{\partial Q}\|^{2}_{F})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(|\frac{\partial x_{i}^{T}QK^{T}x_{j}}{\partial Q}|^{2}_{F})$ (A.31)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|x_{i}x_{j}^{T}K\|^{2}_{F})$ (A.32)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (A.33)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}x_{ik}^{2}(\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (A.34)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}x_{ik}^{2}(\sum_{m=1}^{D}x_{jm}^{2}k_{ml}^{2}+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}x_{jm}k_{ml}x_{jn}k_{nl}))$ (A.35)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}k_{ml}x_{jn}k_{nl}))$ (A.36)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})+0)$ (A.37)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.38)
$=\sum_{i=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\mathop{\mathbb{E}}(x_{ik}^{2}x_{ik}^{2}k_{kl}^{2})+\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\sum_{m=1,m\neq k}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.39)
$=NDd3\sigma_{x}^{4}\sigma_{w}^{2}+N(N-1)D(D-1)d\sigma_{x}^{4}\sigma_{w}^{2}$ (A.40)
$\approx N^{2}D^{2}d\sigma_{x}^{4}\sigma_{w}^{2}.$ (A.41)
@@
$\mathop{\mathbb{E}}(\|\frac{\partial(XQK^{T}X^{T})^{p}}{\partial Q}\|^{2}_{F})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|\frac{(\partial x_{i}^{T}QK^{T}x_{j})^{p}}{\partial Q}\|^{2}_{F})$ (A.42)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|p(x_{i}^{T}QK^{T}x_{j})^{p-1}\frac{\partial x_{i}^{T}QK^{T}x_{j}}{\partial Q}\|^{2}_{F})$ (A.43)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|p(x_{i}^{T}QK^{T}x_{j})^{p-1}x_{i}x_{j}^{T}K\|^{2}_{F})$ (A.44)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(p^{2}(x_{i}^{T}QK^{T}x_{j})^{2p-2}\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2}).$ (A.45)
@@
$\mathop{\mathbb{E}}(\|\frac{\partial(XQK^{T}X^{T})^{p}}{\partial Q}\|^{2}_{F})$ $=$ $p^{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(z_{ij}^{2p-2}\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (A.50)
$=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.51)
$+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}k_{ml}x_{jn}k_{nl}))$ (A.52)
$=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.53)
$+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-3}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2}x_{jn}^{2}k_{nl}^{2}))$ (A.54)
$\approx$ $N^{2}Dd(D(D^{2p-2}d^{p-1}(2p-3)!!\sigma_{x}^{4p}\sigma_{w}^{4p-2})+0$ (A.55)
$=$ $N^{2}D^{2p}d^{p}(2p-3)!!\sigma_{x}^{4p}\sigma_{w}^{4p-2}$ (A.56)
@@ A.2 Experiments
A.2.1 Ablations
In this section we carry out ablations on the experiments we did in the main paper.
Scale ablations:
The theory we developed in section 4.2 suggested that we needed to scale our polynomial activations by $\frac{1}{\sqrt{N}}$ to obtain a complexity bound of $\mathcal{O}(\sqrt{N})$. In general, we could also scale the polynomial activations by $\mathcal{O}(\frac{1}{\sqrt{N}})$ to see if we can get a better accuracr. Figure 10 carries out an ablation on both the ViT-Base architecture and the ViT-Small architecture to see how different scales affect the accuracy for the $x^{3}$ activation. We found that in general a scale from $\frac{1}{8}$ to $\frac{1}{25}$ seemed to perform very well.
Figure 10: We show how the top-1% accuracy changes as the scale of the activation $x^{3}$ for a ViT-Base (left) and ViT-Small (right) architecture. The x-axis plots the denominator of the scale for easier readability. In other words as we go to the right of the x-axis the scale used on the activation $x^{3}$ is getting smaller.
Activation ablations:
Our theory primarily compared the activations $\frac{1}{\sqrt{N}}x^{3}$ and $\frac{1}{\sqrt{N}}x$ with softmax where $N$ was the sequence length of the input. In this section we carry out further experimental comparisons of our activations with other activations. We compare with a variety of activations used in the literature as well as the exponential function $Exp=e^{-x}$ which can be thought of as softmax without the row normalization scaling. Table 5 shows the results of the experiments.
Activation ViT-Base ViT-small
softmax 79.6 80.2
$\frac{x^{3}}{14}$ 79.6 80.5
$\frac{x}{14}$ 76.9 77.8
$\frac{x^{2}}{14}$ 78.7 79.9
$ReLU$ 77.3 77.5
$ELU$ 78.2 78.5
$GELU$ 78.2 78.4
$Tanh$ 77.1 77.3
$Exp$ 77.9 78.0
Table 5: Top-1% accuracy with various activations on a ViT-Base and ViT-Small architecture.
A.2.2 Frobenius norm computations
@@
In section 5.1 we showed plots of the Frobenius norm of the self-attention matrix and for the Jacobian of the self-attention matrix for different scalings of the $x^{3}$ and $x$ activations along with softmax. This was done for a ViT-Tiny architecture on the Tiny-ImageNet dataset. Figure 11 shows the plots of the Frobenius norm of the self-attention matrix for the Tiny-ViT architecture, during training, for all layers averaged over the heads within each layer. Figure 12 shows the Frobenius norm of the Jacobian of the self-attention matrix during training for each layer, averaged over the total number of heads within each layer.
Figures 13 and 14 show similar results for the ViT-Small architecture.
@@
Figure 11: Frobenius norm of self-attention matrix for different scaled activations on ViT-Tiny during training (zoom in for better viewing).
@@
Figure 12: Frobenius norm of Jacobian of self-attention matrix for different scaled activations on ViT-Tiny during training (zoom in for better viewing).
Figure 13: Frobenius norm of self-attention matrix for different scaled activations on ViT-Small during training (zoom in for better viewing).
Figure 14: Frobenius norm of Jacobian of self-attention matrix for different scaled activations on ViT-Small during training (zoom in for better viewing).

### v2

@@ A.1 Theoretical analysis
$\mathop{\mathbb{E}}(c_{11}^{2})$ $\approx\mathop{\mathbb{E}}(\sum_{\{i_{1},\dots,i_{p}\}\in\{1,\dots,D\}}\prod_{k=1}^{p}a_{1,i_{k}}^{2}b_{i_{k},1}^{2})$ (43)
@@
$\mathop{\mathbb{E}}(\|\frac{\partial XQK^{T}X^{T}}{\partial Q}\|^{2}_{F})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(|\frac{\partial x_{i}^{T}QK^{T}x_{j}}{\partial Q}|^{2}_{F})$ (44)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|x_{i}x_{j}^{T}K\|^{2}_{F})$ (45)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (46)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}x_{ik}^{2}(\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (47)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}x_{ik}^{2}(\sum_{m=1}^{D}x_{jm}^{2}k_{ml}^{2}+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}x_{jm}k_{ml}x_{jn}k_{nl}))$ (48)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (49)
$\hskip 28.45274pt+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}k_{ml}x_{jn}k_{nl}))$ (50)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})+0)$ (51)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (52)
$=\sum_{i=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\mathop{\mathbb{E}}(x_{ik}^{2}x_{ik}^{2}k_{kl}^{2})$ (53)
$\hskip 28.45274pt+\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\sum_{m=1,m\neq k}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (54)
$=NDd3\sigma_{x}^{4}\sigma_{w}^{2}+N(N-1)D(D-1)d\sigma_{x}^{4}\sigma_{w}^{2}$ (55)
$\approx N^{2}D^{2}d\sigma_{x}^{4}\sigma_{w}^{2}.$ (56)
@@
$\mathop{\mathbb{E}}(\|\frac{\partial(XQK^{T}X^{T})^{p}}{\partial Q}\|^{2}_{F})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|\frac{(\partial x_{i}^{T}QK^{T}x_{j})^{p}}{\partial Q}\|^{2}_{F})$ (57)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|p(x_{i}^{T}QK^{T}x_{j})^{p-1}\frac{\partial x_{i}^{T}QK^{T}x_{j}}{\partial Q}\|^{2}_{F})$ (58)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|p(x_{i}^{T}QK^{T}x_{j})^{p-1}x_{i}x_{j}^{T}K\|^{2}_{F})$ (59)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(p^{2}(x_{i}^{T}QK^{T}x_{j})^{2p-2}\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2}).$ (60)
@@
$\mathop{\mathbb{E}}(\|\frac{\partial(XQK^{T}X^{T})^{p}}{\partial Q}\|^{2}_{F})$ $=p^{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(z_{ij}^{2p-2}\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (65)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (66)
$+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}k_{ml}x_{jn}k_{nl}))$ (67)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (68)
$+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-3}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2}x_{jn}^{2}k_{nl}^{2}))$ (69)
$\approx N^{2}Dd(D(D^{2p-2}d^{p-1}(2p-3)!!\sigma_{x}^{4p}\sigma_{w}^{4p-2})+0$ (70)
$=N^{2}D^{2p}d^{p}(2p-3)!!\sigma_{x}^{4p}\sigma_{w}^{4p-2}$ (71)
@@ A.2 Experiments
A.2.1 Hardware
The vision transformer experiments in Section 6.1, the object detection and instance segmentation experiments in Section 6.2 and the Nyströmformer experiments from Section 6.3 were all carried out on Nvidia A100 GPUs. The physics informed transformer experiments in Section A.2.5 were carried out on a Nvidia A6000 GPU.
A.2.2 Experimental hyperparameters
Vision transformers in Section 6.1.
In Section 6.1 we tested four different vision transformers, ViT-B [10], DeiT-B [25], Swin-B [17] and XCiT-M [3], with the activations softmax, $x^{3}$ + fixed scaling, $x^{3}$ + dynamic scaling and $x^{3}$. The training strategy follow the exact strategy used in each of the original papers, we used the Timm [28] libraries to train our models.
Object detection and instance segmentation in Section 6.2.
Each of the models are trained for 36 epochs using the AdamW optimizer with learning rate of $10^{-4}$, $0.05$ weight decay and batch size of $16$.
Nyströmformer in Section 6.3.
For the Nyströmformer experiments we used the exact same training strategy form the original paper [30]. Model were trained using the GitHub provided by the authors of [30].
A.2.3 Frobenius norm computations
@@
In Section 5 we showed plots of the Frobenius norm of the self-attention matrix and for the Jacobian of the self-attention matrix for softmax, $\frac{1}{14}x^{3}$, and $x^{3}$. This was done for a ViT-Tiny architecture on the Tiny-ImageNet dataset. Figure 6 shows the plots of the Frobenius norm of the self-attention matrix for the ViT-Tiny architecture, during training, for all layers averaged over the heads within each layer. Figure 7 shows the Frobenius norm of the Jacobian of the self-attention matrix during training for each layer, averaged over the total number of heads within each layer.
@@
Figure 6: Frobenius norm of self-attention matrix for softmax, $\frac{1}{16}x^{3}$ and $x^{3}$ on ViT-Tiny during training on the Tiny-ImageNet dataset (zoom in for better viewing).
@@
Figure 7: Frobenius norm of the Jacobian of the self-attention matrix for softmax, $\frac{1}{16}x^{3}$ and $x^{3}$ on ViT-Tiny during training on the Tiny-ImageNet dataset (zoom in for better viewing).
A.2.4 Discussion on Linear attention and other polynomials on ViTs
In the experiments Section 6 we showed our insights using the polynomial $x^{3}$ as this was a non-linear polynomial that did not satisfy the key properties that softmax did. In this section we will compare softmax with other polynomials that we also found to perform well.
The results for ViTs is shown in Table 5. As we can see from that table each polynomial does far better when scaled using the theory from Section 4. We note that when we tried to train polynomial higher than order $6$ they did not train well. On inspecting the weights we found that several were very close to zero leading to a gradient vanishing problem. We hypothesize that this is because the transformers were randomly initialized with weights about zero. Thus taking such a weight and applying a polynomial of the form $x^{k}$ with $k\geq 6$ would make those weights orders of magnitude smaller, making it difficult to train after some point.
Table 5: Comparisons of pre-training models with different activation functions on ImageNet-1k. We report top-1 classification accuracy (%).
Models
ViT-B DeiT-B Swin-B XciT-M
softmax 80.3 81.5 83.5 82.7
$x$ + fixed scale 78.4 79.4 80.6 80.2
$x$ + dynamic scale 78.7 79.5 80.7 80.4
$x$ 74.1 77.9 78.3 78.1
$x^{2}$ + fixed scale 80.1 81.5 83.4 82.5
$x^{2}$ + dynamic scale 80.3 81.6 83.5 82.7
$x^{2}$ 77.8 78.2 79.8 79.4
$x^{4}$ + fixed scale 80.3 81.5 83.7 82.7
$x^{4}$ + dynamic scale 80.3 81.6 83.7 82.8
$x^{4}$ 77.9 78.6 79.9 79.6
$x^{5}$ + fixed scale 80.3 81.4 83.4 82.5
$x^{5}$ + dynamic scale 80.3 81.5 83.4 82.6
$x^{5}$ 77.7 78.0 79.4 79.5
A.2.5 Physics informed transformers
In [31] transformers for modeling partial differential equations (PDEs) were introduced and shown to yield superior performance on a variety of PDEs. In this section we compare softmax with the polynomial activation we have been using in the previous section on two common PDEs within the literature, namely the convection PDE and the 1d-Reaction PDE.
Convection PDE:
This is a 1-dimensional PDE that models transport phenomena. The PDE system is defined by the following equations
$\frac{\partial u}{\partial t}+\beta\frac{\partial u}{\partial x}=0,\text{ for }x\in[0,2\pi],t\in[0,1]$ (72)
$IC:u(x,0)=sin(x),BC:u(0,t)=u(2\pi,t)$ (73)
where $\beta$ is the convection coefficient and is set as $\beta=50$. IC stands for initial condition and BC for boundary condition. For more details on this system consult [31].
1d-reaction PDE:
The 1d-reaction equation is commonly used in modeling chemical reactions. It is defined by the following system.
$\frac{\partial u}{\partial t}-\rho u(1-u)=0\text{ for }x\in[0,2\pi],t\in[0,1]$ (74)
$IC:u(x,0)=exp\big{(}\frac{(x-\pi)^{2}}{2(\pi/4)^{2}}\big{)},BC:u(0,t)=u(2\pi,t)$ (75)
where $\rho$ is known as the reaction coefficient and is set as $\rho=5$. For more details on this system consult [31].
As in [31] we considered a physics informed transformer (PINNsformer) on the above two PDEs. We compared softmax with $x^{3}$ + fixed scaling, $x^{3}$ + dynamic scaling, where the scale was initialized to $\frac{1}{\sqrt{51}}$ as the input sequence length is $51$, and $x^{3}$. As in [31] we computed the training loss, and for testing the relative mean absolute error (rMAE), and the relative mean squared error (rMSE). Table 6 shows the results from which it can be seen that both $x^{3}$ + fixed scaling and $x^{3}$ + dynamic scaling are competitive with the original softmax.
Table 6: Comparison of different activations on PINNsformer for Convection and 1D-Reaction problems.
Convection PDE 1D-Reaction PDE
Loss rMAE rMSE Loss rMAE rMSE
softmax 5.4e-5 4.1e-2 4.4e-2 3.8e-6 3.1e-2 5.8e-2
$x^{3}$ + fixed scale 4.9e-5 3.9e-2 4.3e-2 3.1e-6 3.0e-2 5.6e-2
$x^{3}$ + dynamic scale 4.1e-5 3.5e-2 4.0e-2 2.8e-6 2.7e-2 5.4e-2
$x^{3}$ 7.2e-5 9.8e-2 1.1e-1 6.2e-6 6.1e-2 9.8e-2
The training of the physics informed transformers followed the exact recipe in the original paper [31] with their data and training schemes available from their GitHub [1]. The evaluation metrics are defined by
$\text{rMAE}=\frac{\sum_{n=1}^{N}\left|\hat{u}(x_{n},t_{n})-u(x_{n},t_{n})\right|}{\sum_{n=1}^{N}\left|u(x_{n},t_{n})\right|}$ (76)
$\text{rMSE}=\sqrt{\frac{\sum_{n=1}^{N}\left|\hat{u}(x_{n},t_{n})-u(x_{n},t_{n})\right|^{2}}{\sum_{n=1}^{N}\left|u(x_{n},t_{n})\right|^{2}}}$ (77)
where $N$ is the number of testing points, $\hat{u}$ is the neural network approximation, and $u$ is the ground truth.